Which wavelength of light has the highest frequency? 14 mm 14 nm 8 mm 8 nm Request Answer Submit
Answer #1

For highest frequency, wavelength should be smallest as wavelength and frequency are inversely related.

Option 1:

Wavelength = 14 mm

= 1.4*10^-2 m

Option 2:

Wavelength = 14 nm

= 1.4*10^-8 m

Option 3:

Wavelength = 8 mm

= 8*10^-3 m

Option 4:

Wavelength = 8 nm

= 8*10^-9 m

Clearly wavelength is smallest for option 4

So, frequency is highest in this case

Answer: option 4